my house and the school are 16 miles apart. The scale on a map is 1.5 in =4 miles how far apart should the two cities be
Answer by stanbon(75887)

my house and the school are 16 miles apart. The scale on a map is 1.5 in =4 miles how far apart should the two cities be on the map.
--------
Ans:: (16/4)*1.5" = 4*1.5" = 6"
----------
Cheers,
Stan H.